How many different committees of 5people can be chosen from 10 people

1
Expert's answer
2022-06-13T16:27:42-0400

Solution:

Let's give each people one number: 1, 2, 3, 4, 5, 6, 7, 8, 9, 10;

We can write it like set: "A={1,2,3,4,5,6,7,8,9,10}", this is set with 10 elements. Now, we need to find number of subsets of set A which equal number of different committees with 5 people:

"N=\frac{n!}{(n-m)!m!}" - this is formula of number of subsets with m elements in set with n elements.

"N=\frac{10!}{(10-5)!\times5!}=\frac{10!}{5!\times5!}=\frac{6\times7\times8\times9\times10}{5!}=252;"

Answer:

"N=252" -number of different committees with 5 people.